# Think Then Embed: Generative Context Improves Multimodal Embedding

### v2

@@ 1 Introduction
Multimodal embedding-based retrieval has emerged as a popular and effective solution for handling diverse data types such as text, images, and videos (Radford et al., 2021; Li et al., 2023a; Yu et al., 2022; Li et al., 2023b). Traditionally, these models focus on learning general-purpose representations that capture content similarity across modalities. Recently, there has been growing interest in instruction-aware Universal Multimodal Embeddings (UME) (Jiang et al., 2024; Gu et al., 2025). UMEs are powerful representations because they bridge the gap between general-purpose content similarity and user-specific requirements. This capability unlocks a wide range of applications, such as retrieving documents or images based on nuanced queries, powering personalized recommendation systems, supporting multimodal search with complex transformations, and tackling knowledge-intensive tasks where the same input may yield different embeddings depending on the downstream objective. To advance this direction, recent benchmarks like MMEB (Meng et al., 2025) aggregate diverse, instruction-driven retrieval tasks—including VQA, grounding, and document retrieval—where both the input and the user instruction determine the retrieval target.
In this context, leveraging multimodal large language models (MLLMs) as encoders for UME is a promising new approach (Jiang et al., 2024; Gu et al., 2025; Lan et al., 2025). MLLMs can generate instruction-aware representations that incorporate long context and nuanced user guidance (Peng et al., 2024). However, while state-of-the-art models on these benchmarks have advanced through techniques such as hard negative sampling (Lee et al., 2025; Thirukovalluru et al., 2025; Lan et al., 2025; Lin et al., 2025), additional training stages (Gu et al., 2025; Yu et al., 2025; Lin et al., 2025), external data (Yu et al., 2025; Chen et al., 2025b), and improved embedding extraction (Faysse et al., 2024), a central challenge remains: understanding instructions of varying complexity that require different levels of reasoning. Addressing this challenge calls for leveraging the generative capacity of MLLMs, rather than restricting them to encoder-only models.
@@
Our approach is motivated by a closer examination of the tasks in MMEB-v2 (Meng et al., 2025), revealing that many, such as VQA, classification, visual grounding, and composed retrieval, require substantial reasoning capabilities. For example, the first case in Fig. 1 (from the MMEB-v2 RefCOCO dataset) involves retrieving the image region referred to by the query “vehicle second closest to camera”. Rather than simply encoding the query and matching it to the target image, we argue that it is beneficial for the model to first reason about the query—identifying, for instance, that it refers to the vehicle with “bright yellow on the upper half and deep blue on the lower”. For such complex tasks, we believe that explicitly augmenting multi-step reasoning and contextual grounding enables the model to achieve more fine-grained and accurate retrieval.
To address these challenges, we propose Think-Then-Embed (TTE) for UME. The key idea is to introduce an explicit thinking stage before embedding, in which the model generates reasoning traces based on the given instruction. Prior work has shown that intermediate reasoning processes, such as Chain-of-Thought (CoT) (Wei et al., 2023b), can significantly improve the accuracy of language model generation. We hypothesize that a similar benefit can be achieved for multimodal representation learning, since the embeddings produced by MLLMs are inherently conditioned on the sequence of preceding tokens. In this work, we investigate how explicit chain-of-thought reasoning can enhance universal multimodal embeddings, by enabling the model to better interpret and follow complex task instructions.
In this paper, we made several contributions. First, we proposed a Think-Then-Embed (TTE) framework that by exploiting reasoning traces generated by a powerful reasoner (e.g. Qwen2.5 VL 72B), it can achieve state-of-the-art performance on MMEB V2 (Meng et al., 2025) with a smaller embedder (Qwen2 7B), surpassing close-sourced models trained with additional data. This demonstrates that CoT can also benefit representation learning.
@@
Finally, to improve efficiency of both inference and number of parameters in the TTE framework, we explore the capabilities of instruction reasoning and embedding creation within a single unified model. Although prior work (Yu et al., 2025) also investigated the unification of generation and contrastive learning, it treats generation and embedding as separate, unrelated tasks. In contrast, our focus is on enhancing embedding through reasoning, that is, generate first, then embed. We focus on two strategies for this unification: (1) joint contrastive-autoregressive training of the same backbone for reasoning and embedding tasks, and (2) a two-stage autoregressive-then-contrastive training with a dedicated embedding head on top of the reasoner. Empirically, we find that the two-stage approach consistently outperforms joint training, almost halving the overall parameters, while not degrading end-to-end TTE retrieval performance.
@@ 2 Related Work
Multimodal Embeddings. The multimodal representation paradigm has been popularized by large-scale foundational models such as CLIP (Radford et al., 2021; Xu et al., 2024), BLIP (Li et al., 2023a), and SigLIP (Zhai et al., 2023), which encode images and texts using separate uni-modality encoders and enforce alignment using contrastive objectives.
Universal Multimodal Embedding. Recently there has been growing interest in developing Universal Multimodal Embedding (UME) (Gu et al., 2025; Jiang et al., 2024; Meng et al., 2025), where the embedding depends both on the query and task instruction. Representative examples include VLM2Vec (Jiang et al., 2024), which proposes the Massive Multimodal Embedding Benchmark (MMEB), comprising a wide range of cross / multimodal retrieval tasks, as well as non-conventional retrieval tasks such as VQA, grounding, and classification. Subsequently, MMEB-V2 (Meng et al., 2025) is proposed, extending MMEB-V1 to include video and visual document (visdoc) tasks.
MLLM-based Embedding Models. Recent studies have gone beyond dual encoder setups, building embedding models directly on top of powerful Multimodal Language Models (MLLMs) (Jiang et al., 2024; Meng et al., 2025; Gu et al., 2025; Yu et al., 2025; Thirukovalluru et al., 2025). Existing work on MLLM-based embedding typically explores a training strategy such as text-continual contrastive training (Gu et al., 2025; Yu et al., 2025; Chen et al., 2025a), hard negative mining approach (Lee et al., 2025; Lan et al., 2025; Thirukovalluru et al., 2025), additional training data (Zhou et al., 2024; Chen et al., 2025b), and architectural design (e.g. with / without causal mask) (Lee et al., 2025; Chen et al., 2025a). However, these methods all treat MLLMs solely as encoders, while overlooking their generative capability obtained from pre-training. In contrast, we explore an orthogonal approach, where we leverage MLLMs for both generative reasoning and representation learning.
LLM-based Query Rewriting. Frequently used in text-based retrievals, query rewriting (Ma et al., 2023) is an effective approach to improve retrieval accuracy. Popular trends include prompt-based rewriting (Wilson et al., 2025; Ye et al., 2023), iterative or Reinforcement Learning (RL)-guided (Cao et al., 2025; Zhu et al., 2025) query rewritings, and multi-reformulation approaches (Kostric and Balog, 2024; Dhole et al., 2024). However, these works all focus on text-based retrievals using external retrievers, where no learned embedding is involved.
Recently, Bai et al. (2025b) applied the idea of query rewriting to uni-modal encoder-based models (e.g. CLIP) for text-to-video retrieval, by enriching textual queries with additional context to bridge information asymmetry. In contrast, our work focuses on leveraging reasoning to improve the quality of general instruction-following multimodal embeddings, which can apply to both query and target.
@@ 3 Preliminary: MLLM-based Universal Multimodal Embedding
where $f_{\theta}$ denotes the MLLM embedder and Pooling refers to the pooling operation for aggregating MLLM’s hidden states into the final embedding. Following previous works (Jiang et al., 2024; Gu et al., 2025), we use the last token’s hidden state for final embedding. During training, we adopt the standard uni-directional $q\rightarrow t$ InfoNCE loss as follow:
@@ 4 Think-then-Embed
We propose our Think-then-Embed (TTE) framework for enhancing Universal Multimodal Embeddings (UME) by incorporating an explicit reasoning step prior to the encoding step. TTE consists of a reasoner, that generates an Embedding-Centric Reasoning (ECR) trace, and an embedder, that produces task-specific representations conditioned on both the original input and the reasoning trace generated by the reasoner. Figure 2 compares the standard approach of directly encoding multimodal inputs for embedding (2(a)) with our proposed approach TTE (2(b), 2(c)).
@@ 4.1 Embedding-Centric Reasoning
Here, we introduce the concept of Embedding-Centric Reasoning (ECR), a form of intermediate reasoning tailored to improve embedding quality. Inspired by chain-of-thought (CoT) reasoning (Wei et al., 2023b) in complex problem solving, ECR are generative reasoning traces that explicitly support the production of target embeddings. For example, for VQA query, the ECR represents the model’s step-by-step reasoning to understand the query, whereas for grounding query, it captures a detailed description of the referred object along with its surrounding visual context.
@@ 4.3 TTE with Finetuned Reasoner
Encouraged by these results, we further finetune a dedicated reasoner, initializing it from the same backbone as the embedder (2B or 7B). We finetune the small reasoner using ECR traces generated by the teacher reasoner (e.g., Qwen2.5-72B). Given a multimodal input $\langle\mathcal{V},[\texttt{Ins}],\mathcal{T}\rangle$, we maximize the conditional likelihood of the ECR $\psi$ by optimizing the standard negative log-likelihood (NLL) loss in LLM autoregressive next prediction finetuning:
@@
Figure 3: Embedding head designs: (a) Attention Pooling with learnable query. (b) NV-Embed-style (Lee et al., 2025) pooler. (c) Qformer-style embedding head, and (d) Embedding head with self attention blocks initialized from the backbone MLLM. Green denotes trainable components. Q, K and V denote query, key, and value in attention mechanism. MHSA refers to MultiHead Self Attention, and   refers to output embedding.
@@ 4.4 TTE with Unified Reasoner and Embedder
A Systematic Study of Pluggable Embedding Head. We present the first systematic study of embedding head designs that can be seamlessly plugged into a frozen backbone MLLM. This stands in contrast to the focus of prior LLM-based embedding models (Lee et al., 2025; Lin et al., 2025; Jiang et al., 2024), which primarily rely on either full finetuning or parameter-efficient approaches based on LoRA (Hu et al., 2021). As illustrated in Fig. 3, we evaluate four embedding head designs: (1) Learnable attention queries over backbone hidden states (Fig. 3(a)); (2) Learnable latent context with backbone hidden states as queries; (3) A QFormer-style (Li et al., 2023a) embedding head patched onto the last $n$ layers of the backbone (Fig. 3(c)). (4) Repeat the last $n$ layers of the backbone to formulate N new layers as the embedding head (Fig. 3(d)).
@@ 5.1 Experimental Setup
Datasets. We conduct our experiments on both MMEB-V2 (Meng et al., 2025) and MMEB-V1 (Jiang et al., 2024) datasets. MMEB-V1 comprises 20 in-distribution (IND) tasks and 16 out-of-distribution (OOD) tasks, spanning 4 meta-tasks: classification, VQA, retrieval, and grounding. MMEB-V2 extends MMEB-V1 by introducing additional video and visual document (visdoc) retrieval tasks. Video tasks including video QA, classification, retrieval, and video moment retrieval. Visdoc involves supportive document retrieval, where the task is to retrieve supporting documents given the query. Overall, MMEB-V2 includes a total of 78 test tasks. We report NDCG@5 for visdoc retrieval and Precision@1 for image and video tasks (Meng et al., 2025).
Models. We employ Qwen2-VL 2B and 7B as the backbone, and introduce three variants of TTE: TTE with a large teacher reasoner (TTEt), TTE with a small, SFT-ed student reasoner (TTEs), and TTE with a unified reasoner and embedder (TTEu). We compare our approach against several baselines, including dual-encoder methods such as CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), UniIR (Wei et al., 2023a), and MagicLens (Zhang et al., 2024). We also directly compare with VLM2Vec-V1 (Jiang et al., 2024) and VLM2Vec-V2 (Meng et al., 2025) with Qwen2-VL as the backbone MLLM. Additionally, we evaluate against recent MLLM-based embedding approaches, including UniME (Gu et al., 2025), LLaVE (Lan et al., 2025), and B3 (Thirukovalluru et al., 2025).
Training Details. For both MMEB-V1 and MMEB-V2, we use a global batch size of $8192$ , with a learning rate of $2e^{-4}$ for the backbone MLLM and $5e^{-4}$ for the embedding head. The temperature for contrastive loss is set to $0.02$. We train the backbone using LoRA (Hu et al., 2021) with rank $16$ and alpha $64$. Following VLM2Vec, we employ GradCache (Gao et al., 2021) to increase the per-device batch size. We train for $1$ epoch on MMEB-V1, and $2.3$ epochs for MMEB-V2. For MMEB-V2, we follow the same data weighing setup as in VLM2Vec-V2, and adopt its interleaved sampling strategy, where one global batch is splitted into $n$ sub batches, each from one dataset.
For supervised finetuning of the ECR Reasoner, we perform full parameter finetuning of the MLLM while keeping the visual encoder frozen, using a learning rate of $2e^{-5}$, a global batch size of $128$, and training for an epoch. We use DeepSpeed (Aminabadi et al., 2022) for optimizer offloading.
@@ 6.1 Main results
CLIP (Radford et al., 2021) - 42.8 9.1 53.0 51.8 37.1 38.7 37.8
SigLIP (Zhai et al., 2023) - 40.3 8.4 31.6 59.5 32.3 38.0 34.8
UniIR ($\text{CLIP}_{\text{SF}}$) (Wei et al., 2023a) - 44.3 16.2 61.8 65.3 47.1 41.7 44.7
MagicLens (Zhang et al., 2024) - 38.8 8.3 35.4 26.0 31.0 23.7 27.8
@@
VLM2Vec-V1 (Jiang et al., 2024) Qwen2VL 59.0 49.4 65.4 73.4 66.0 52.6 59.3
UniME (Gu et al., 2025) LLaVA-1.6 54.8 55.9 64.5 81.8 68.2 52.7 64.2
LLaVE (Lan et al., 2025) Aquila-VL 62.1 60.2 65.2 84.9 69.4 59.8 65.2
B3 (Thirukovalluru et al., 2025) Qwen2VL 67.0 61.2 70.9 79.9 72.1 63.1 68.1
@@
VLM2Vec-V1 (Jiang et al., 2024) Qwen2VL 62.6 57.8 69.9 81.7 65.2 56.3 65.8
UniME (Gu et al., 2025) LLaVA-OV 66.8 66.6 70.5 90.9 74.6 65.8 70.7
LLaVE (Lan et al., 2025) LLaVA-OV 65.7 65.4 70.9 91.9 75.0 64.4 70.3
B3 (Thirukovalluru et al., 2025) Qwen2VL 70.0 66.5 74.1 84.6 75.9 67.1 72.0
QQMM (Xue et al., 2025) LLaVA-OV 69.9 70.0 72.1 86.0 77.2 66.6 72.5
@@
Table 1: Results on the MMEB-V1 benchmark (Jiang et al., 2024). The scores are averaged per meta-task. Best/2nd-best performance across both 2B and 7B categories are in bold/underlined.
@@
Table 2: Results on the MMEB-V2 benchmark (Meng et al., 2025). Best/2nd-best performance across all models are in bold/underlined. Task abbreviations: CLS (classification), QA (question answering), RET (retrieval), GD (grounding), MRET (moment retrieval), VDR (ViDoRe), VR (VisRAG), and OOD (out-of-domain).

### v4

@@ 1 Introduction
Multimodal embedding-based retrieval has emerged as a popular and effective solution for handling diverse data types such as text, images, and videos (clip; blip2; coca; flip). Traditionally, these models focus on learning general-purpose representations that capture content similarity across modalities. Recently, there has been growing interest in instruction-aware Universal Multimodal Embeddings (UME) (vlm2vec; unime). UMEs are powerful representations because they bridge the gap between general-purpose content similarity and user-specific requirements. This capability unlocks a wide range of applications, such as retrieving documents or images based on nuanced queries, powering personalized recommendation systems, supporting multimodal search with complex transformations, and tackling knowledge-intensive tasks where the same input may yield different embeddings depending on the downstream objective. To advance this direction, recent benchmarks like MMEB (vlm2vecv2) aggregate diverse, instruction-driven retrieval tasks—including VQA, grounding, and document retrieval—where both the input and the user instruction determine the retrieval target.
In this context, leveraging multimodal large language models (MLLMs) as encoders for UME is a promising new approach (vlm2vec; unime; llave). MLLMs can generate instruction-aware representations that incorporate long context and nuanced user guidance (peng2024answer). However, while state-of-the-art models on these benchmarks have advanced through techniques such as hard negative sampling (nv_embed; b3; llave; mm_embed), additional training stages (unime; cafe; mm_embed), external data (cafe; mmE5), and improved embedding extraction (fayssecolpali), a central challenge remains: understanding instructions of varying complexity that require different levels of reasoning. Addressing this challenge calls for leveraging the generative capacity of MLLMs, rather than restricting them to encoder-only models.
@@
Our approach is motivated by a closer examination of the tasks in MMEB-v2 (vlm2vecv2), revealing that many, such as VQA, classification, visual grounding, and composed retrieval, require substantial reasoning capabilities. For example, the first case in Fig. 1 (from the MMEB-v2 RefCOCO dataset) involves retrieving the image region referred to by the query “vehicle second closest to camera”. Rather than simply encoding the query and matching it to the target image, we argue that it is beneficial for the model to first reason about the query—identifying, for instance, that it refers to the vehicle with “bright yellow on the upper half and deep blue on the lower”. For such complex tasks, we believe that explicitly augmenting multi-step reasoning and contextual grounding enables the model to achieve more fine-grained and accurate retrieval.
To address these challenges, we propose Think-Then-Embed (TTE) for UME. The key idea is to introduce an explicit thinking stage before embedding, in which the model generates reasoning traces based on the given instruction. Prior work has shown that intermediate reasoning processes, such as Chain-of-Thought (CoT) (cot), can significantly improve the accuracy of language model generation. We hypothesize that a similar benefit can be achieved for multimodal representation learning, since the embeddings produced by MLLMs are inherently conditioned on the sequence of preceding tokens. In this work, we investigate how explicit chain-of-thought reasoning can enhance universal multimodal embeddings, by enabling the model to better interpret and follow complex task instructions.
In this paper, we made several contributions. First, we proposed a Think-Then-Embed (TTE) framework that by exploiting reasoning traces generated by a powerful reasoner (e.g. Qwen2.5 VL 72B), it can achieve state-of-the-art performance on MMEB V2 (vlm2vecv2) with a smaller embedder (Qwen2 7B), surpassing close-sourced models trained with additional data. This demonstrates that CoT can also benefit representation learning.
@@
Finally, to improve efficiency of both inference and number of parameters in the TTE framework, we explore the capabilities of instruction reasoning and embedding creation within a single unified model. Although prior work (cafe) also investigated the unification of generation and contrastive learning, it treats generation and embedding as separate, unrelated tasks. In contrast, our focus is on enhancing embedding through reasoning, that is, generate first, then embed. We focus on two strategies for this unification: (1) joint contrastive-autoregressive training of the same backbone for reasoning and embedding tasks, and (2) a two-stage autoregressive-then-contrastive training with a dedicated embedding head on top of the reasoner. Empirically, we find that the two-stage approach consistently outperforms joint training, almost halving the overall parameters, while not degrading end-to-end TTE retrieval performance.
@@ 2 Related Work
Multimodal Embeddings. The multimodal representation paradigm has been popularized by large-scale foundational models such as CLIP (clip; metaclip), BLIP (blip2), and SigLIP (siglip), which encode images and texts using separate uni-modality encoders and enforce alignment using contrastive objectives.
Universal Multimodal Embedding. Recently there has been growing interest in developing Universal Multimodal Embedding (UME) (unime; vlm2vec; vlm2vecv2), where the embedding depends both on the query and task instruction. Representative examples include VLM2Vec (vlm2vec), which proposes the Massive Multimodal Embedding Benchmark (MMEB), comprising a wide range of cross / multimodal retrieval tasks, as well as non-conventional retrieval tasks such as VQA, grounding, and classification. Subsequently, MMEB-V2 (vlm2vecv2) is proposed, extending MMEB-V1 to include video and visual document (visdoc) tasks.
MLLM-based Embedding Models. Recent studies have gone beyond dual encoder setups, building embedding models directly on top of powerful Multimodal Language Models (MLLMs) (vlm2vec; vlm2vecv2; unime; cafe; b3). Existing work on MLLM-based embedding typically explores a training strategy such as text-continual contrastive training (unime; cafe; moca), hard negative mining approach (nv_embed; llave; b3), additional training data (megapairs; mmE5), and architectural design (e.g. with / without causal mask) (nv_embed; moca). However, these methods all treat MLLMs solely as encoders, while overlooking their generative capability obtained from pre-training. In contrast, we explore an orthogonal approach, where we leverage MLLMs for both generative reasoning and representation learning.
LLM-based Query Rewriting. Frequently used in text-based retrievals, query rewriting (query_rewriting) is an effective approach to improve retrieval accuracy. Popular trends include prompt-based rewriting (wilson2025contextualizingsearchqueriesincontext; ye2023enhancingconversationalsearchlarge), iterative or Reinforcement Learning (RL)-guided (cao2025icriterativeclarificationrewriting; zhu2025convsearchr1enhancingqueryreformulation) query rewritings, and multi-reformulation approaches (kostric2024surprisinglysimpleeffectivemultiquery; dhole2024generativequeryreformulationusing). However, these works all focus on text-based retrievals using external retrievers, where no learned embedding is involved.
Recently, bai2025bridginginformationasymmetrytextvideo applied the idea of query rewriting to uni-modal encoder-based models (e.g. CLIP) for text-to-video retrieval, by enriching textual queries with additional context to bridge information asymmetry. In contrast, our work focuses on leveraging reasoning to improve the quality of general instruction-following multimodal embeddings, which can apply to both query and target.
@@ 3 Preliminary: MLLM-based Universal Multimodal Embedding
where $f_{\theta}$ denotes the MLLM embedder and Pooling refers to the pooling operation for aggregating MLLM’s hidden states into the final embedding. Following previous works (vlm2vec; unime), we use the last token’s hidden state for final embedding. During training, we adopt the standard uni-directional $q\rightarrow t$ InfoNCE loss as follow:
@@ 4 Think-then-Embed
We propose our Think-then-Embed (TTE) framework for enhancing Universal Multimodal Embeddings (UME) by incorporating an explicit reasoning step prior to the encoding step. TTE consists of a reasoner, that generates an Embedding-Centric Reasoning (ECR) trace, and an embedder, that produces task-specific representations conditioned on both the original input and the reasoning trace generated by the reasoner. Fig. 2 compares the standard approach of directly encoding multimodal inputs for embedding (2(a)) with our proposed approach TTE (2(b), 2(c)).
@@ 4.1 Embedding-Centric Reasoning
Here, we introduce the concept of Embedding-Centric Reasoning (ECR), a form of intermediate reasoning tailored to improve embedding quality. Inspired by chain-of-thought (CoT) reasoning (cot) in complex problem solving, ECR are generative reasoning traces that explicitly support the production of target embeddings. For example, for VQA query, the ECR represents the model’s step-by-step reasoning to understand the query, whereas for grounding query, it captures a detailed description of the referred object along with its surrounding visual context.
@@ 4.3 TTE with Finetuned Reasoner
Encouraged by these results, we further finetune a dedicated reasoner, initializing it from the same backbone as the embedder (2B or 7B). We finetune the small reasoner using ECR traces generated by the teacher reasoner (e.g., Qwen2.5-72B). Given a multimodal input $\langle\mathcal{V},[\texttt{Ins}],\mathcal{T}\rangle$, we maximize the conditional likelihood of the ECR $\psi$ by optimizing the standard negative log-likelihood (NLL) loss in standard LLM autoregressive finetuning:
@@
Figure 3: Embedding head designs: (a) Attention Pooling with learnable query. (b) NV-Embed-style (nv_embed) pooler. (c) Qformer-style embedding head, and (d) Embedding head with self attention blocks initialized from the backbone MLLM. Green denotes trainable components. Q, K and V denote query, key, and value in attention mechanism. MHSA refers to MultiHead Self Attention, and   refers to output embedding.
@@ 4.4 TTE with Unified Reasoner and Embedder
A Systematic Study of Pluggable Embedding Head. We present the first systematic study of embedding head designs that can be seamlessly plugged into a frozen backbone MLLM. This stands in contrast to the focus of prior LLM-based embedding models (nv_embed; mm_embed; vlm2vec), which primarily rely on either full finetuning or parameter-efficient approaches based on LoRA (lora). As illustrated in Fig. 3, we evaluate four embedding head designs: (1) Learnable attention queries over backbone hidden states (Fig. 3(a)); (2) Learnable latent context with backbone hidden states as queries; (3) A QFormer-style (blip2) embedding head patched onto the last $n$ layers of the backbone (Fig. 3(c)). (4) Repeat the last $n$ layers of the backbone to formulate N new layers as the embedding head (Fig. 3(d)).
@@ 5.1 Experimental Setup
Datasets. We conduct our experiments on both MMEB-V2 (vlm2vecv2) and MMEB-V1 (vlm2vec) datasets. MMEB-V1 comprises 20 in-distribution (IND) tasks and 16 out-of-distribution (OOD) tasks, spanning 4 meta-tasks: classification, VQA, retrieval, and grounding. MMEB-V2 extends MMEB-V1 by introducing additional video and visual document (visdoc) retrieval tasks. Video tasks including video QA, classification, retrieval, and video moment retrieval. Visdoc involves supportive document retrieval, where the task is to retrieve supporting documents given the query. Overall, MMEB-V2 includes a total of 78 test tasks. We report NDCG@5 for visdoc retrieval and Precision@1 for image and video tasks (vlm2vecv2).
Models. We employ Qwen2-VL 2B and 7B as the backbone, and introduce three variants of TTE: TTE with a large teacher reasoner (TTEt), TTE with a small, SFT-ed student reasoner (TTEs), and TTE with a unified reasoner and embedder (TTEu). We compare our approach against several baselines, including dual-encoder methods such as CLIP (clip), SigLIP (siglip), UniIR (uniir), and MagicLens (magiclens). We also directly compare with VLM2Vec-V1 (vlm2vec) and VLM2Vec-V2 (vlm2vecv2) with Qwen2-VL as the backbone MLLM. Additionally, we evaluate against recent MLLM-based embedding approaches, including UniME (unime), LLaVE (llave), and B3 (b3).
Training Details. For both MMEB-V1 and MMEB-V2, we use a global batch size of $8192$ , with a learning rate of $2e^{-4}$ for the backbone MLLM and $5e^{-4}$ for the embedding head. The temperature for contrastive loss is set to $0.02$. We train the backbone using LoRA (lora) with rank $16$ and alpha $64$. Following VLM2Vec, we employ GradCache (gradcache) to increase the per-device batch size. We train for $1$ epoch on MMEB-V1, and $2.3$ epochs for MMEB-V2. For MMEB-V2, we follow the same data weighing setup as in VLM2Vec-V2, and adopt its interleaved sampling strategy, where one global batch is splitted into $n$ sub batches, each from one dataset.
For supervised finetuning of the ECR Reasoner, we perform full parameter finetuning of the MLLM while keeping the visual encoder frozen, using a learning rate of $2e^{-5}$, a global batch size of $128$, and training for an epoch. We use DeepSpeed (deepspeed) for optimizer offloading.
@@ 6.1 Main results
CLIP (clip) - 42.8 9.1 53.0 51.8 37.1 38.7 37.8
SigLIP (siglip) - 40.3 8.4 31.6 59.5 32.3 38.0 34.8
UniIR ($\text{CLIP}_{\text{SF}}$) (uniir) - 44.3 16.2 61.8 65.3 47.1 41.7 44.7
MagicLens (magiclens) - 38.8 8.3 35.4 26.0 31.0 23.7 27.8
@@
VLM2Vec-V1 (vlm2vec) Qwen2VL 59.0 49.4 65.4 73.4 66.0 52.6 59.3
UniME (unime) LLaVA-1.6 54.8 55.9 64.5 81.8 68.2 52.7 64.2
LLaVE (llave) Aquila-VL 62.1 60.2 65.2 84.9 69.4 59.8 65.2
B3 (b3) Qwen2VL 67.0 61.2 70.9 79.9 72.1 63.1 68.1
@@
VLM2Vec-V1 (vlm2vec) Qwen2VL 62.6 57.8 69.9 81.7 65.2 56.3 65.8
UniME (unime) LLaVA-OV 66.8 66.6 70.5 90.9 74.6 65.8 70.7
LLaVE (llave) LLaVA-OV 65.7 65.4 70.9 91.9 75.0 64.4 70.3
B3 (b3) Qwen2VL 70.0 66.5 74.1 84.6 75.9 67.1 72.0
QQMM (qqmm) LLaVA-OV 69.9 70.0 72.1 86.0 77.2 66.6 72.5
@@
Table 1: Results on the MMEB-V1 benchmark (vlm2vec). The scores are averaged per meta-task. Best/2nd-best performance across both 2B and 7B categories are in bold/underlined.
@@
Table 2: Results on the MMEB-V2 benchmark (vlm2vecv2). Best/2nd-best performance across all models are in bold/underlined. Task abbreviations: CLS (classification), QA (question answering), RET (retrieval), GD (grounding), MRET (moment retrieval), VDR (ViDoRe), VR (VisRAG), and OOD (out-of-domain).